# Entropic Risk Optimization in Discounted MDPs: Sample Complexity Bounds with a Generative Model

### v2

@@ 2.2 Entropic Risk Preferences
The entropic risk preferences is rooted in expected utility theory. Consider for $\beta\neq 0$ the class of utility functions $u(t)=\frac{1}{\beta}(1-e^{-\beta t})$ defined for $t\in\mathbb{R}$. The utility $u$ is supposed to describe the preferences of some economic agent in the form of how much utility $u(t)$ she derives from some monetary quantity $t\in\mathbb{R}$. For any bounded random variable $X\in L^{\infty}(\Omega,\mathcal{F},\mathbb{P})$, the certainty equivalent $u^{-1}(\mathbb{E}[u(X)])=\frac{-1}{\beta}\log(\mathbb{E}[e^{-\beta X}])$ expresses the amount of money that would give the same utility as that of entering in the bet given by the random variable $X$. We thus define the functional $\rho:L^{\infty}(\Omega,\mathcal{F},\mathbb{P})\rightarrow\mathbb{R}$ by
$\rho(X)=-\frac{1}{\beta}\log(\mathbb{E}[e^{-\beta X}])\,.$
We note that there does not seem to be consensus on wether the functional is parametrized with $\beta$ of $\beta$. We follow this convention considering its widespread use in the actuarial literature asienkiewicz2017note , but we remark that some other lines of work appear to prefer the other parametrization. For the case $\beta\rightarrow 0$, we recover the risk-neutral case that is simply the expectation: $\lim_{\beta\to 0}\rho(X)=\mathbb{E}[X]$. It is straightforward to see that $\rho$ admits the following properties:
@@
$\rho(X)\geq\mathbb{E}[X]\quad\text{for }\beta<0,$ (4)
where properties (3)-(4) follow from Jensen’s inequality. Using $\rho$ as a measure of the preference for different random variables, it follows directly from (2)-(4) that $\rho(X)\leq\rho(\mathbb{E}[X])$ for $\beta>0$ and that $\rho(X)\geq\rho(\mathbb{E}[X])$ for $\beta<0$. If further shows that $\beta>0$ is associated with risk-aversion, while $\beta<0$ is associated with risk-seeking behavior.
Applying risk in a stochastic dynamic process can be done in several ways and is thus more complicated than for a single-period problem. Two approaches to this end exist in the literature. The first and most intuitive one, which is often called the static or non-recursive approach, is to apply the functional to the total discounted sum of rewards $\rho\big(\sum_{t=0}^{\infty}\gamma^{t}r_{t}\big)$, which is well-defined under the bounded rewards assumption, i.e., $r_{t}\in[0,1]$ for all $t$. The other approach, which we may call the recursive approach, works by applying the risk at every step $t$ where we give the details below. The non-recursive approach is probably most intuitive but has several drawbacks. Even though there is no obvious optimality equation for the non-recursive case, a solution to the planning problem has been proposed in hau2023entropic . In comparison, the planning problem is more straightforward with recursive risk-preferences due to the availability of an optimality equation that allows for simple value-iteration type algorithms. The recursive approach also guarantees the existence of an optimal stationary deterministic policy whereas with non-recursive risk preferences the optimal policy might not be time-consistent (see jaquette1976utility ). In this paper, we study the problem with recursive risk-preferences.
@@ 2.3 Value Function and Q-function
A bit of notation is required in order to define the state value function $V$ (henceforth $V$-function) and state-action value function $Q$ (henceforth $Q$-function) of a policy.
@@
We follow the approach of asienkiewicz2017note and bauerle_markov_2024 but since none of their cases include our $\beta>0$ case and also only cover $V$-functions, we give in Section C the full setup with history-dependent policies as well as a full definition of the $V$ and $Q$-functions, and prove existence of a stationary optimal policy and show that the value functions of this policy satisfy a Bellman optimality equation and similarly that value functions of any policy satisfy a Bellman recursion relation. We give an outline here that only deals with stationary policies, which is justified by the results of Section C.
Let $v\in\mathbb{R}^{S}$ and $\pi:\mathcal{S}\rightarrow\mathcal{A}$ be a stationary deterministic policy. We then define $\rho_{s,a}:\mathbb{R}^{S}\rightarrow\mathbb{R}$ by
$\rho_{s,a}(v)=-\frac{1}{\beta}\log\bigg(\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[e^{-\beta v(s^{\prime})}]\bigg)$ (5)
@@
and slighltly abusing the notation, we write $\rho_{s,\pi}$ when $a=\pi(s)$, that is $\rho_{s,\pi}:=\rho_{s,\pi(s)}$. We then define the operator $J_{\pi}:\mathbb{R}^{S}\rightarrow\mathbb{R}^{S}$ given by $J_{\pi}(v)(s)=r(s,\pi(s))+\gamma\rho_{s,\pi}(v)$. The $N$-step total discounted utility $J_{N}(s,\pi)$ is defined as applying $J_{\pi}$ recursively $N$-times to the 0-map, that is $J_{N}(s,\pi):=J_{\pi}^{N}(\boldsymbol{0})(s)$. Note that the outer-most iteration corresponds to the immediate time-step. By properties (1)-(2) of $\rho$, it follows that $J_{N}(s,\pi)$ is increasing in $N$ and that $J_{N}(s,\pi)\leq\frac{1}{1-\gamma}$, so that the limit $N\rightarrow\infty$ exists. This limit is considered the value of state $s$ under the policy $\pi$ of the agent: $V^{\pi}(s)=\lim_{N\rightarrow\infty}J_{N}(s,\pi)$.
@@
The problem of the agent is then for all initial states $s\in\mathcal{S}$ to find a policy $\pi^{*}$ that solves $J(s,\pi^{*})=\sup_{\pi}J(s,\pi)$. In asienkiewicz2017note , the authors consider a more general framework that is not restricted to finite MDPs or stationary policies; they prove that under some conditions —that are trivially fulfilled in the case of finite MDPs— there exists a stationary policy $\pi^{*}$ that maximizes the state-value function for all states $s\in\mathcal{S}$ simultaneously. We show that any optimal policy $\pi^{*}$ that solves the agent satisfies the optimality equation:
@@
$V^{*}(s)=\max_{a\in\mathcal{A}}\bigg(R(s,a)-\frac{\gamma}{\beta}\log\bigg(\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[e^{-\beta V^{*}(s^{\prime})}]\bigg)\bigg)\,$, (6)
@@
where $V^{*}$ is the optimal $V$-function. Also for any stationary deterministic policy $\pi$, we have the Bellman recursion:
@@
$V^{\pi}(s)=R(s,\pi(s))-\frac{\gamma}{\beta}\log\bigg(\mathbb{E}_{s^{\prime}\sim P(\cdot|s,\pi(s))}[e^{-\beta V^{\pi}(s^{\prime})}]\bigg)\,.$ (7)
@@
Since we are not interested only in the planning problem but also in learning, we also introduce the state-action value function $Q$. The approach is very similar to that of the value-function. Given $\pi$, we define the operator $L_{\pi}:\mathbb{R}^{S}\rightarrow\mathbb{R}^{S\times A}$ as follows: for all $v:\mathcal{S}\to\mathbb{R}$, $L_{\pi}(v)(s,a)=R(s,\pi(s))+\gamma\rho_{s,\pi}(v)\,.$ We define the operator $L:\mathbb{R}^{S}\rightarrow\mathbb{R}^{S\times A}$ for all $v:\mathcal{S}\rightarrow\mathbb{R}$ as: $L(v)(s,a)=R(s,a)+\gamma\rho_{s,a}(v)\,.$ We define the $N$-step total discounted utility of the state-action pair $(s,a)$ under $\pi$ as $L_{N}(s,a,\pi):=(L\circ J_{\pi}^{N-1}(\boldsymbol{0}))(s,a)$ and the limit is denoted $Q^{\pi}(s,a)$: $Q^{\pi}(s,a)=\lim_{N\rightarrow\infty}L_{N}(s,a,\pi)$. Although the authors do not consider state-action value functions in their paper, repeating the arguments of asienkiewicz2017note it suffices to consider stationary policies when wanting to solve $\max_{\pi}Q^{\pi}(s,a)$ for all $(s,a)$ and that the solution $Q^{*}$ solves the optimality equation:
$Q^{*}(s,a)=R(s,a)-\frac{\gamma}{\beta}\log\bigg(\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[e^{-\beta\max_{a^{\prime}}Q^{*}(s^{\prime},a^{\prime})}]\bigg)\,.$ (8)
Similarly, it is clear that the $Q$-functions satisfy the Bellman recursion relations:
@@
$Q^{\pi}(s,a)=R(s,a)-\frac{\gamma}{\beta}\log\big(\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[e^{-\beta V^{\pi}(s^{\prime})}]\big)\,.$ (9)
@@ 2.4 Learning Performance
We consider two types of RL algorithms $\mathcal{U}$, namely those that output a $Q$-function $Q_{T}^{\mathcal{U}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ and those that output a policy $\pi_{T}^{\mathcal{U}}:\mathcal{S}\rightarrow\mathcal{A}$ using $T$ transition samples. Note that any algorithm that outputs a $Q$-function also outputs a policy, namely the one obtained by acting greedily with respect to the $Q$-function. There are also ways to obtain a $Q$-function from a policy. There is however no canonical way to do this as the algorithm cannot simply output $Q^{\pi^{\mathcal{U}}}$ since the algorithm does not have access to the true MDP. The way we evaluate the quality of an algorithm that outputs a $Q$-function is by $\|Q^{*}-Q_{T}^{\mathcal{U}}\|$. For an algorithm that instead outputs a policy, we evaluate the policy in terms of $\|V^{*}-V^{\pi_{T}^{\mathcal{U}}}\|$. Often we will suppress $T$ from the notation.
Definition 1 ($(\varepsilon,\delta)$-correctness).
We say that an algorithm $\mathcal{U}$ that outputs a $Q$-function $Q^{\mathcal{U}}$ is $(\varepsilon,\delta)$-correct on a set of MDPs $\mathbb{M}$ if $\mathbb{P}(\|Q^{*}-Q^{\mathcal{U}}\|\leq\varepsilon)\geq 1-\delta$ for all $M\in\mathbb{M}$. Similarly, we say that an algorithm $\mathcal{U}$ that outputs a policy $\pi^{\mathcal{U}}$ is $(\varepsilon,\delta)$-correct on a set of MDPs $\mathbb{M}$ if $\mathbb{P}(\|V^{*}-V^{\mathcal{\pi^{\mathcal{U}}}}\|\leq\varepsilon)\geq 1-\delta$ for all $M\in\mathbb{M}$.
3 Model-Based Risk-Sensitive Q-Value Iteration
In this section we describe the model-based value iteration algorithm which aims at finding the optimal $Q$-function $Q^{*}$. We then give an upper bound on the total number of calls to the generative model needed in order for this algorithm to be $(\varepsilon,\delta)$-correct. The model based approach is based on working on an MDP, which may disagree with the true MDP because it does not use the true transition probabilities but an estimate of the transition functions obtained from $n$ calls to each of the state-action pairs in $\mathcal{S}\times\mathcal{A}$ as described in the algorithm below.

### v3

@@ 3.1 Entropic Risk Preferences
The entropic risk measure (ERM) is rooted in expected utility theory [54]. Consider for $\beta\neq 0$ the class of utility functions $u(t)=\frac{1}{\beta}(1-e^{-\beta t})$ defined for $t\in\mathbb{R}$. The utility $u$ is supposed to describe the preferences of some economic agent in the form of how much utility $u(t)$ she derives from some monetary quantity $t\in\mathbb{R}$. For any bounded random variable $X\in L^{\infty}(\Omega,\mathcal{F},\mathbb{P})$, it is easy to verify that the associated certainty equivalent to $u$ is $u^{-1}(\mathbb{E}[u(X)])=\frac{-1}{\beta}\log(\mathbb{E}[e^{-\beta X}])$, which expresses the amount of money that would give the same utility as that of entering in the bet given by the random variable $X$. We thus define the functional $\rho:L^{\infty}(\Omega,\mathcal{F},\mathbb{P})\rightarrow\mathbb{R}$ by [^4] We note that there is a lack of consensus regarding the sign of $\beta$ in the definition of ERM. We follow this convention considering its widespread use in the actuarial literature [4]. We refer to Appendix A for a related discussion.
$\rho(X):=\rho(X;\beta):=-\frac{1}{\beta}\log\big(\mathbb{E}[e^{-\beta X}]\big)\,.$ (1)
Evidently, when $\beta\rightarrow 0$ we recover the risk-neutral case, which simply coincides with the expectation: $\lim_{\beta\to 0}\rho(X)=\mathbb{E}[X]$. Further, it is straightforward to see that $\rho$ admits the following:
@@
$\rho(X)\geq\mathbb{E}[X],\quad\text{for }\beta<0,$ (5)
where properties (4)-(5) follow from Jensen’s inequality. Using $\rho$ as a measure of the preference for different random variables, it follows directly from (3)-(5) that $\rho(X)\leq\rho(\mathbb{E}[X])$ for $\beta>0$ and that $\rho(X)\geq\rho(\mathbb{E}[X])$ for $\beta<0$. It further shows that $\beta>0$ is associated with risk-aversion, while $\beta<0$ is associated with a risk-seeking behavior. It is well-known that the ERM, unlike CVaR, does not belong to the nice class of coherent risk measures; we refer the reader to Appendix A for a primer on risk measures, where we collect some definitions and concrete examples.
3.2 Discounted Markov Decision Processes with Entropic Risk
We write the 6-tuple $M=(\mathcal{S},\mathcal{A},P,R,\gamma,\beta)$ to denote a finite, discounted infinite-horizon Markov decision process (MDP), where $\mathcal{S}=\{1,2,\ldots,S\}$ is the finite state space of size $S:=|\mathcal{S}|$, $\mathcal{A}=\{1,2,\ldots,A\}$ is the finite action space of size $A:=|\mathcal{A}|$, $P:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the transition probability function, $R:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is the reward function, $\gamma\in(0,1)$ is the discount factor, and $\beta\neq 0$ is the risk-parameter. We use $Z=\mathcal{S}\times\mathcal{A}$ to denote the set of all state-action pairs, and write $P_{s,a}(s^{\prime})$ as short-hand notation for $P(s^{\prime}|s,a)$ for any $(s,a)\in Z$. For simplicity of exposition, we consider a deterministic reward function, as is standard in the literature. The agent interacts with the MDP $M$ as follows. At the beginning of the process, $M$ is in some initial state $s_{0}\in\mathcal{S}$. At each time $t\geq 0$, the agent is in state $s_{t}\in\mathcal{S}$ and decides on an action $a_{t}\in\mathcal{A}$ according to some rule. The MDP generates a reward $r_{t}:=R(s_{t},a_{t})$ and a next-state $s_{t+1}\sim P(\cdot|s_{t},a_{t})$. The MDP moves to $s_{t+1}$ when the next time slot begins, and this process continues ad infinitum. This process yields a growing sequence $(s_{t},a_{t},r_{t})_{t\geq 0}$. The agent’s goal is to maximize an objective function, as a function of the collected rewards $(r_{t})_{t\geq 0}$, which depends on both $\gamma$ and $\beta$.
To concretely define the agent’s objective using ERM, we discuss two approaches of applying the functional $\rho$ in (1) in the context of MDPs. The first approach, called non-recursive (or static or non-iterated) [32, 52], consists in applying $\rho$ to the total discounted sum of rewards $\rho\big(\sum_{t=0}^{\infty}\gamma^{t}r_{t}\big)$, which is well-defined under the bounded rewards assumption, i.e., $r_{t}\in[0,1]$ for all $t$. This problem admits no obvious optimality equation, although solution and approximation schemes for planning have been proposed in the literature [32, 52]. The other approach where $\rho$ is applied at every step is called recursive (also called dynamic or iterated) [4]. The planning problem in this case is tractable thanks to existence of Bellman-type optimality equations. The recursive approach also guarantees the existence of an optimal stationary deterministic policy, whereas the non-recursive approach may lead to optimal policies that are not time-consistent (see [36]). In this paper, we study the case where the RL objective is defined via the recursive ERM.
@@ 3.3 Value Function and Q-function
We shall introduce some notation and definitions to formally define the value function $V$ and state-action value function $Q$ (henceforth, $Q$-value) of a policy. We follow the approach of [4] and [6], but since none of their cases include our $\beta<0$ case and also only cover value function, we present in Appendix B the full setup with history-dependent policies as well as a thorough definition of the value and $Q$-values. There, we prove existence of a stationary optimal policy, and show that the value functions of this policy satisfy a Bellman optimality equation and that the value of any policy satisfies a Bellman recursion. We give an outline here that only deals with stationary policies, which is justified by the results of Appendix B.
Let $v\in\mathbb{R}^{S}$ and $\pi:\mathcal{S}\rightarrow\mathcal{A}$ be a stationary deterministic policy. We define $\rho_{s,a}:\mathbb{R}^{S}\rightarrow\mathbb{R}$ as
$\rho_{s,a}(v)=-\frac{1}{\beta}\log\Big(\mathbb{E}_{s^{\prime}\sim P_{s,a}}[e^{-\beta v(s^{\prime})}]\Big)$ (6)
@@
and slightly abusing the notation, we write $\rho_{s,\pi}$ when $a=\pi(s)$, i.e., $\rho_{s,\pi}:=\rho_{s,\pi(s)}$. We then introduce the operator $J_{\pi}:\mathbb{R}^{S}\rightarrow\mathbb{R}^{S}$ with $J_{\pi}(v)(s)=R(s,\pi(s))+\gamma\rho_{s,\pi}(v)$. The $N$-step total discounted utility $J_{N}(s,\pi)$ is defined as applying $J_{\pi}$ recursively $N$ times to the 0-map: $J_{N}(s,\pi):=J_{\pi}^{N}(\boldsymbol{0})(s)$. Note that the outer-most iteration corresponds to the immediate time-step. Then, the value of policy $\pi$ is defined as: $V^{\pi}(s)=\lim_{N\rightarrow\infty}J_{N}(s,\pi)$ for all $s\in\mathcal{S}$. By properties (2)-(3), it follows that $J_{N}(s,\pi)$ is increasing in $N$ and that $J_{N}(s,\pi)\leq\frac{1}{1-\gamma}$ for all $s\in\mathcal{S}$, so that the limit above exists and the value function is thus well-defined. The optimal state-values are defined as $V^{*}(s)=\sup_{\pi}V^{\pi}(s)$ for all $s\in\mathcal{S}$, where the $\sup$ is taken over all possible policies. Any policy achieving $V^{*}(s)$ at all states is called optimal and $V^{*}$ is called the optimal value function. Further, given $\varepsilon>0$, a policy achieving $V^{\pi}(s)\geq V^{*}(s)-\varepsilon$ for all $s\in\mathcal{S}$ is called $\varepsilon$-optimal.
@@
In [4], the authors consider a more general MDP framework that is not restricted to finite MDPs or stationary policies; they prove that under some conditions —that are trivially fulfilled in the case of finite MDPs— there exists a stationary deterministic optimal policy in the $\beta>0$ case. This result is easily extended to the $\beta<0$ case and we give a unified proof for completeness in Appendix B. The optimal value function $V^{*}$ satisfies the Bellman optimality equation:
@@
$V^{*}(s)=\max_{a\in\mathcal{A}}\bigg(R(s,a)\!-\!\frac{\gamma}{\beta}\log\bigg(\mathbb{E}_{s^{\prime}\sim P_{s,a}}[e^{-\beta V^{*}(s^{\prime})}]\bigg)\bigg)\,,\quad\forall s\in\mathcal{S}.$
@@
Further, for any stationary deterministic policy $\pi$, the value function satisfies the Bellman recursion:
@@
$V^{\pi}(s)=R(s,\pi(s))-\frac{\gamma}{\beta}\log\bigg(\mathbb{E}_{s^{\prime}\sim P_{s,\pi(s)}}[e^{-\beta V^{\pi}(s^{\prime})}]\bigg)\,,\quad\forall s\in\mathcal{S}.$ (7)
@@
We introduce the $Q$-value functions using a similar approach. Given $\pi:\mathcal{S}\to\mathcal{A}$, we define the operator $L_{\pi}:\mathbb{R}^{S}\rightarrow\mathbb{R}^{S\times A}$ as follows: for all $v:\mathcal{S}\to\mathbb{R}$, for all $(s,a)$, $L_{\pi}(v)(s,a)=R(s,\pi(s))+\gamma\rho_{s,\pi}(v)\,.$ Also, we define the operator $L:\mathbb{R}^{S}\rightarrow\mathbb{R}^{S\times A}$ as follows: for all $v:\mathcal{S}\rightarrow\mathbb{R}$, for all $(s,a)$, $L(v)(s,a)=R(s,a)+\gamma\rho_{s,a}(v)\,.$ We define the $N$-step total discounted utility of the state-action pair $(s,a)$ under $\pi$ as $L_{N}(s,a,\pi):=(L\circ J_{\pi}^{N-1}(\boldsymbol{0}))(s,a)$ and the limit is denoted $Q^{\pi}(s,a)$: $Q^{\pi}(s,a)=\lim_{N\rightarrow\infty}L_{N}(s,a,\pi)$. Although [4] does not consider Q-value functions, building on their arguments we show in Appendix B that it suffices to consider stationary policies when solving $\max_{\pi}Q^{\pi}(s,a)=:Q^{*}(s,a)$ for all $(s,a)$ and that $Q^{*}$ satisfies the optimality equation:
@@
$Q^{*}(s,a)\!=\!R(s,a)\!-\!\frac{\gamma}{\beta}\log\Big(\mathbb{E}_{s^{\prime}\sim P_{s,a}}\big[e^{-\beta\max_{a^{\prime}}Q^{*}(s^{\prime},a^{\prime})}\big]\Big)\,,\quad\forall(s,a)\in\mathcal{S}\times\mathcal{A}.$
Similarly, the $Q$-value of a policy $\pi$ satisfies the Bellman recursion:
$Q^{\pi}(s,a)=R(s,a)-\frac{\gamma}{\beta}\log\Big(\mathbb{E}_{s^{\prime}\sim P_{s,a}}\big[e^{-\beta Q^{\pi}(s^{\prime},\pi(s^{\prime}))}\big]\Big)\,,\quad\forall(s,a)\in\mathcal{S}\times\mathcal{A}.$
Let us introduce the operators $\mathcal{T}^{\pi},\mathcal{T}:\mathbb{R}^{S\!\times\!A}\rightarrow\mathbb{R}^{S\times A}$, which for $f:\mathcal{S}\!\times\!\mathcal{A}\rightarrow\mathbb{R}$ are defined as
$(\mathcal{T}f)(s,a)$ $=R(s,a)-\frac{\gamma}{\beta}\log\sum_{s^{\prime}}P_{s,a}(s^{\prime})e^{-\beta\max_{a^{\prime}}f(s^{\prime},a^{\prime})}\,$,
$(\mathcal{T}^{\pi}f)(s,a)$ $=R(s,a)-\frac{\gamma}{\beta}\log\sum_{s^{\prime}}P_{s,a}(s^{\prime})e^{-\beta f(s^{\prime},\pi(s^{\prime}))}\,.$
Then, the Bellman equations above can be written as $Q^{*}=\mathcal{T}Q^{*}$ and $Q^{\pi}=\mathcal{T}^{\pi}Q^{\pi}$. [^5] We note that our analysis only rests on the Bellman optimality equation; the Bellman equations for $V^{\pi}$ and $Q^{\pi}$ are included for completeness.
We conclude this section by a remark about the case of rewards bounded in $R_{\min},R_{\max}]$.
Remark 1.
For rewards bounded in $R_{\min},R_{\max}]$, one can equivalently consider rewards in $[0,1]$ but with a risk parameter $\frac{\beta}{w}$, with $w:=R_{\max}-R_{\min}$. This is verified by observing that (i) $\rho$ is translation invariant, implying that the range $w$ is important –not the absolute values– so that one must model rewards as belonging to $[0,w]$; and (ii) one has $\rho(wX,\beta)=-\frac{w}{w\beta}\log(\mathbb{E}[e^{-\beta wX}]=w\rho(X,w\beta)$ for any $w>0$, so that scaling the range by $w$ amounts to working with $\rho$ with a risk parameter $w\beta$. This observation justifies our focus on rewards in $[0,1]$, but more importantly signals that the risk parameter is calibrated to a specific reward range.
@@ 3.4 Learning Performance
We consider RL algorithms that aim to find an $\varepsilon$-optimal policy or an $\varepsilon$-optimal value function for input $\varepsilon>0$ under ERM, while having access to a generative model (or simulator) of the MDP. Precisely speaking, the generative model can produce a sample $s^{\prime}\sim P_{s,a}$ for any queried state-action $(s,a)$. We consider two types of such algorithms, which we generically denote by $\mathcal{U}$: The first type outputs a $Q$-value $Q_{T}^{\mathcal{U}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, whereas the second outputs a policy $\pi_{T}^{\mathcal{U}}:\mathcal{S}\rightarrow\mathcal{A}$ using $T$ samples. We evaluate the quality of an algorithm that outputs a $Q$-value by $\|Q^{*}-Q_{T}^{\mathcal{U}}\|$. For an algorithm that instead outputs a policy, we evaluate it in terms of $\|V^{*}-V^{\pi_{T}^{\mathcal{U}}}\|$. Often, we will suppress $T$ from the notation. This leads to the notion of $(\varepsilon,\delta)$-correct value and policy for input parameters $(\varepsilon,\delta)$:
Definition 1 ($(\varepsilon,\delta)$-correct value and policy).
An algorithm $\mathcal{U}$ that outputs a $Q$-value $Q^{\mathcal{U}}$ is called $(\varepsilon,\delta)$-value-correct on a set of MDPs $\mathbb{M}$ if $\mathbb{P}(\|Q^{*}-Q^{\mathcal{U}}\|\leq\varepsilon)\geq 1-\delta$ for all $M\in\mathbb{M}$. Similarly, an algorithm $\mathcal{U}$ that outputs a policy $\pi^{\mathcal{U}}$ is called $(\varepsilon,\delta)$-policy-correct on a set of MDPs $\mathbb{M}$ if $\mathbb{P}(\|V^{*}-V^{\mathcal{\pi^{\mathcal{U}}}}\|\leq\varepsilon)\geq 1-\delta$ for all $M\in\mathbb{M}$.
The notion of $(\varepsilon,\delta)$-value-correctness yields a sample complexity notion in the case of value learning, while $(\varepsilon,\delta)$-policy-correctness serves a similar role for policy learning.
We remark that any algorithm that outputs a $Q$-value also outputs a policy, e.g., the one obtained by acting greedily with respect to the $Q$-value. However, it is well-known that such a greedy policy can be off by a factor of $1/(1-\gamma)$, which impacts the corresponding sample complexity of policy learning; see discussions in [68, 67, 1]. To remedy this, the literature resort to proof arguments, which may however come at a cost of limiting the $\varepsilon$-range, as briefly discussed in Section 2.
4 Model-Based ERM Q-Value Iteration
In this section, we present a model-based algorithm, called MB-RS-QVI, for value and policy learning settings under ERM, assuming access to a generative model of the MDP. Then, we derive PAC-type bounds on its sample complexity.
We begin with introducing the protocol for obtaining $N$ samples from each state-action pair in $Z=\mathcal{S}\times\mathcal{A}$; this is done by making a total of $T=NSA$ calls to the generative model (see Algorithm 1).